# Uniqueness of Galilean and Carrollian limits of gravitational theories and application to higher derivative gravity

### 01 Jan 2024 v1

@@ I Introduction
The Galilean limit of general relativity (GR) was first computed in [1] as an extension of Newton-Cartan (NC) theory by taking the limit as the speed of light goes to infinity, $c\to\infty$. Later development were made in [2, 3] showing that the leading order (LO) in the Galilean limit gives the Newton-Cartan structure, but the next-to-leading order (NLO) gives the type II torsional Newton Cartan (TNC) geometry [4]. With the growing interest in non-Lorentzian geometry in the recent years, there have been many papers studying this limit [5, 6, 7, 8, 9]. A convenient parametrization for taking such a limit, the pre-nonrelativistic (PNR) parametrization, was introduced in [10] and further used in [11]; for a review see [12]. The Galilean limit can also result from performing the ADM decomposition, then taking the limit of the norm of the normal vector to infinity. We call it the Infinite scaling (IS) method. Another method of constructing NC and TNC theories from Lorentzian ones is by gauging the Bargmann algebra [13]. This was motivated by the fact that the Galilean algebra, which on gauging gives NC geometry, can be constructed by the nonrelativistic Inönü-Wigner contraction of the Poincaré group. On extending it to the Bargmann algebra, the gauging procedure gives TNC geometry. This method, known as Galilean algebra gauging (GAG), was applied to study the Galilean limit of GR in [14]. Since then, the Galilean limit has been used in condensed matter [15, 16], fluid mechanics [17, 18] and string theory [19, 20, 21] where the NC and TNC geometries are promoted to their stringy versions.
@@
As of today, the Carrollian limit has been applied in various areas of theoretical physics. Specifically, the Carrollian physics and Carrollian structures were analyzed in the context of representations of the Carroll group, i.e, the Carroll particles [28, 29, 30, 31]. In condensed matter physics [32, 33, 34], the Carrollian limit was used to study fractons. It also found various applications in field theory [35, 36, 37, 38] and conformal field theory [39, 40, 41, 42], such as the holography. The Carrollian limit of fluids [43, 44, 45, 46, 47, 48] was found particularly interesting as it allowed non trivial motion even in the LO. Some cosmological applications of the limit were also considered [49, 50]. The Carrollian limit was further used in string theory [51, 52, 53], gravity [54, 55, 56, 57, 58, 59, 60, 61, 62], black holes [24, 63, 64, 31, 65]; particularly to analyze dynamics of particles near black-hole horizons [66, 67, 68]. Recently, the Carrollian limit gave rise to a new theory of holography on null boundaries of asymptotically flat spacetimes [69, 70, 40, 71].
@@
  4. iv)
    By rescaling certain terms and taking the appropriate limit to get the correct theory [74] [^1] Since rescaling terms will give a subset of theories given by the PUL approach, we consider it at a special case..
In this paper, we show that the PNR parametrization can be constructed from GAG. Similarly, the PUL parametrization can be constructed from CAG. We also show that the PUL and ZS approaches of taking the Carrollian limit are completely equivalent. Throughout this paper, we consider the ADM decomposition with an extra real parameter $\epsilon$ which can be interpreted as a signature if it is $1$, $-1$, or 00; the zero signature case is a Carrollian manifold [73]. Once we establish the uniqueness of the Galilean and Carrollian limits (i.e., the equivalence of above approaches), we construct an algorithm for computing the $n$-th order of the Galilean, and the Carrollian expansions of a completely generic HDG. The problem is transformed into computationally easier optimization problems which upon solving gives the desired order.
@@
  • In Sec. III, we give a proof that the PUL and the ZS approaches are equivalent in the sense that they describe the same limit of the metric with same variables and parameters. Furthermore, we show that the PUL parametrization and the Carroll compatible connection can be constructed by gauging the Carroll algebra, i.e., the equivalence of PUL and CAG. Similarly, we also prove that we can construct the PNR parametrization and the Galilei compatible connection from the Galilei algebra gauging procedure, i.e., the equivalence of PNR and GAG.
@@ II Review of non-Lorentzian limits
We begin with a brief review of the methods that are used in the available literature for construction non-Lorentzian theories.
@@ II.1 Pre-ultralocal parametrization
The PUL parameterization is the parametrization that is suitable for taking the ultralocal (Carrollian) limit [58]. The metric parametrization is the same as (3) with the conditions (4), but a different connection that is compatible with Carrollian symmetries. This can be achieved by demanding $V^{\mu}$ and $\Pi_{\mu\nu}$ to be covariantly constant,
@@ II.2 Pre-nonrelativistic parametrization
A convenient parametrization of the metric that allows one to take the nonrelativistic (Galilean) limit is the PNR parametrization [62]. It reads
It proves useful to introduce a Galilei compatible connection by demanding that $\Pi^{\mu\nu}$ and $T_{\mu}$ are covariantly constant,
@@
Now, assuming the analyticity of all the quantities, we expand $\Pi_{\mu\nu},\Pi^{\mu\nu},V^{\mu},T_{\mu}$ in powers of $\tfrac{1}{c^{2}}$:
@@ II.3 Zero signature approach
The ZS approach is a method of taking the ultralocal limit of a Lorentzian theory. It consists of performing the ADM decomposition, and then setting the signature of the manifold to zero. This process results in a Carrollian manifold [73]. The first step is to decompose the metric
$g_{\mu\nu}=n_{\mu}n_{\nu}+h_{\mu\nu},\hskip 10.0pt$ $\hskip 10.0ptg^{\mu\nu}=n^{\mu}n^{\nu}+h^{\mu\nu},$ (7)
where $h_{\mu\nu}$ is the induced metric on the spatial surfaces and $n_{\mu}$ is the corresponding orthogonal 1-form. In a Lorentzian manifold we have $n_{\mu}n^{\mu}=-1$. Contrary to the previously reviewed approaches, here we use the usual Levi-Civita connection. Expressing the Lagrangian in terms of the ADM variables, then setting the signature of the manifold to zero, we get a degenerate metric (on a Carrollian manifold) with the Carrollian theory.
II.4 Carroll algebra gauging
@@
The second equality follows from $\xi^{\mu}\omega_{\mu}^{a}=\xi^{\mu}\omega_{\mu\nu}^{ab}=0$. Let $\omega$ be a real parameter. We notice that the rescaling of $\tau_{\mu}$ and $\zeta$ by $\tfrac{1}{\omega}$ and of $\omega_{\mu}{}^{a}$ and $\lambda^{a}$ by $\omega$ leaves the connection and the infinitesimal transformation parameter invariant. The field strength tensor (curvature) associated to $\mathcal{A}_{\mu}$ as
@@
II.5 Galilei algebra gauging
@@
where $G_{a}$ denote the Galilean boosts. The general connection and the infinitesimal transformation parameter are the same as in the Carrollian case. However, the rescaling of $\tau_{\mu}$ and $\zeta$ by $L$, and of $\omega_{\mu}{}^{a}$ and $\lambda^{a}$ by $\tfrac{1}{\omega}$ leaves the connection and the infinitesimal transformation parameter invariant. The curvatures are given by
@@
where $\partial_{\mu}$ is the partial derivative with respect to a generic coordinate system. Taking the limit $L\to\infty$ gives Galilean theories.
@@ III.1 Equivalence of PUL and ZS approaches
Here, we rewrite the formulae used in the ADM decomposition with an extra real parameter $\epsilon$. This parameter should be understood as a modification to the normalization of the vector $n_{\mu}$ that is orthogonal to the spacelike slices, ${n_{\mu}n^{\mu}=\epsilon}$. The metric is Lorentzian for $\epsilon<0$, Riemannian for $\epsilon>0$, and degenerate for $\epsilon=0$. Following the usual ADM decomposition procedure, we write the metric and its inverse in an adapted coordinate system satisfying $\boldsymbol{\partial}_{t}=N\boldsymbol{n}+\boldsymbol{N}$,
@@
$g^{\mu\nu}=\begin{bmatrix}\tfrac{1}{\epsilon N^{2}}&\tfrac{N^{j}}{N^{2}}\\ \tfrac{N^{i}}{N^{2}}&h^{ij}+\epsilon\frac{N^{i}N^{j}}{N^{2}}\end{bmatrix}.$ (17)
@@
It is easy to show that $h_{\mu}^{\nu}h_{\nu}^{\rho}=h^{\rho}_{\mu}$, and $h^{\mu}_{\nu}=\delta_{\nu}^{\mu}-\tfrac{1}{\epsilon}n^{\mu}n_{\nu}$. The extrinsic curvature is then given by
$K_{\mu\nu}=h_{\mu}^{\alpha}h_{\nu}^{\beta}\nabla_{\alpha}n_{\beta}=-\nabla_{\nu}n_{\nu}+\tfrac{1}{\epsilon}n_{\mu}n^{\alpha}\nabla_{\alpha}n_{\nu},$ (20)
It is easy to deduce that $K_{\mu\nu}=-\tfrac{1}{2}\pounds_{\boldsymbol{n}}h_{\mu\nu}$. We notice that although (18) were written in some adapted coordinates the fact that $n_{\mu}$ has an overall factor of $\epsilon$ while $n^{\mu}=O(1)$ as $\epsilon\to 0$ does not change, that is because the dependency on the signature does not depend on the choice of the coordinate system. Thus, we can rescale $n_{\mu}=-\epsilon\tilde{n}_{\mu}$ where $\tilde{n}_{\mu}=O(1)$ when $\epsilon\to 0$. Writing (19) in terms of the covector $\tilde{n}_{\mu}$ we get
@@
Having shown that, we write Gauss relation and its contractions in terms of the parameter $\epsilon$. Beginning with the Gauss relation:
@@
where $K=h^{\mu\nu}K_{\mu\nu}$, $R_{\sigma\nu}$ is the 4-dimensional Ricci tensor and $\bar{R}_{\alpha\beta}$ is the 3-dimensional Ricci tensor on a spacelike slice. Now contracting $\alpha$ and $\beta$, we get
@@ III.2 Equivalence of PNR and IS approaches
Similar to the previous subsection, beginning with (21). Taking the limit $\epsilon\to\infty$ results in the metric in [62] in the PNR parameterization. Thus, all the curvature tensors in both approaches must be equivalent. All the formulae afterwards are valid for the Galilean limit provided we take the limit $\epsilon\to\infty$.
@@ III.3 Equivalence of PUL and CAG approaches
Let us now reconstruct the PUL parametrization from the CAG procedure. Using the procedure delineated in II.4, we define inverses to the gauge fields vielbein ($\tau_{\mu}$ and $e_{\mu}{}^{a}$) by
@@
Performing the redefinitions in [14], where $\tau_{\mu}$ , $\omega_{\mu}{}^{a}$, $\zeta$, and $\lambda^{a}$ are rescaled by $\tfrac{1}{\omega}$ and taking the limit $\omega\to\infty$, leaving $\mathcal{A}_{\mu}$ and $\Pi$ invariant, we can write the metric and its inverse as
@@
This form is the same form in the PUL approach with $\omega=\tfrac{1}{c^{2}}$. It follows from this fact is that all curvature tensors will be the same in the two approaches. Thus, any gravity theory will be expanded similarly, i.e, the two approaches are equivalent.
We now derive the Carroll compatible connection presented in [58] from the CAG approach. To do that, define a covariant derivative which are compatible with the veilbein [14] by
@@
For more general connections see [56]. This procedure is the generalized version of the one in appendix B of [58]. It is easy to see from this point using (13) that the torsion and the curvature tensors match that of the PUL ones, for example the Riemann tensor is given by
@@
which, upon calculating explicitly, it gives the same form as eq.(2.18) in [77].

### 23 Jan 2024 v2

@@ I Introduction
The Galilean limit of general relativity (GR) was first computed in [1] as an extension of Newton-Cartan (NC) theory by taking the limit as the speed of light goes to infinity, $c\to\infty$. Later development were made in [2, 3] showing that the leading order (LO) in the Galilean limit gives the Newton-Cartan structure, but the next-to-leading order (NLO) gives the type II torsional Newton-Cartan (TNC) geometry [4]. With the growing interest in non-Lorentzian geometry in the recent years, there have been many papers studying this limit [5, 6, 7, 8, 9]. The methods of taking the Galilean limit of a relativistic theory are
  1. i)
    By using the pre-nonrelativistic (PNR) parametrization of the metric which was introduced in [10] and further used in [11]; for a review see [12].
  2. ii)
    By performing the ADM decomposition, then taking the limit of the norm of the normal vector to infinity. We call it the “infinite scaling” (IS) method.
  3. iii)
    By gauging the Bargmann algebra [13]. This is motivated by the fact that the Galilean algebra, which on gauging gives NC geometry, can be constructed by the nonrelativistic Inönü-Wigner contraction of the Poincaré group. On extending it to the Bargmann algebra, the gauging procedure gives TNC geometry. This method, known as Galilean algebra gauging (GAG), was applied to study the Galilean limit of GR in [14].
A different method was proposed in [15] where covariant and contravariant tensors are scaled differently (by $c^{-1}$ and $c$ respectively), although this method makes it easier to take the Galilean limits of Maxwell’s theory to the LO and NLO, it is equivalent to the PNR approach since lowering and raising indices is done by the metric with a factor of $c^{2}$ and its inverse with a factor of $c^{-2}$, and the two approaches give the same Galilean theories. The localization of Galilean symmetries are done in [16, 17, 18]. Since then, the Galilean limit has been used in condensed matter [19, 20], fluid mechanics [21, 22] and string theory [23, 24, 25] where the NC and TNC geometries are promoted to their stringy versions.
@@
As of today, the Carrollian limit has been applied in various areas of theoretical physics. Specifically, the Carrollian physics and Carrollian structures were analyzed in the context of representations of the Carroll group, i.e, the Carroll particles [32, 33, 34, 35]. In condensed matter physics [36, 37, 38], the Carrollian limit was used to study fractons and its continuum field limit [39, 40]. It also found various applications in field theory [41, 42, 43, 44] and conformal field theory [45, 46, 47, 48], such as the holography. The Carrollian limit of fluids [49, 50, 51, 52, 53, 54] was found particularly interesting as it allowed non trivial motion even in the LO. Some cosmological applications of the limit were also considered [55, 56]. The Carrollian limit was further used in string theory [57, 58, 59], gravity [60, 61, 62, 63, 64, 65, 66, 67, 68], black holes [28, 69, 70, 35, 71]; particularly to analyze dynamics of particles near black-hole horizons [72, 73, 74]. Recently, the Carrollian limit gave rise to a new theory of holography on null boundaries of asymptotically flat spacetimes [75, 76, 46, 77].
@@
Another possibility is by rescaling certain terms and taking the appropriate limit to get the desired theory [80]. This method was implemented for multigravity theories in [81]. Since rescaling terms will give a subset of theories given by the PUL approach, we consider it at a special case. The Carrollian limit of GR was also constructed using the Kol-Smolkin (KS) decomposition (which is dual to the ADM decomposition), expanding the resulting quantities in powers of $c^{-1}$, then taking the limit ${c\to 0}$ [82]. This method gives the same results as the ZS approach since the two decompositions are dual, and the limit ${c\to 0}$ gives the same result as setting the signature to zero as we will show.
In this paper, we show that the PNR parametrization can be constructed from GAG. Similarly, the PUL parametrization can be constructed from CAG. We also show that the PUL and ZS approaches of taking the Carrollian limit are completely equivalent (and, similarly, PNR and IS). Throughout this paper, we consider the ADM decomposition with an extra real parameter $\epsilon$ which can be interpreted as a signature if it is $1$, $-1$, or 00; the zero signature case is a Carrollian manifold [79]. Once we establish the uniqueness of the Galilean and Carrollian limits (i.e., the equivalence of above approaches), we construct an algorithm for computing the $n$-th order of the Galilean, and the Carrollian expansions of a completely generic higher derivative gravity (HDG). The problem is transformed into computationally easier optimization problems which upon solving gives the desired order.
@@
  • In Sec. III, we give a proof that the PUL and the ZS approaches are equivalent in the sense that they describe the same limit of the metric with same variables and parameters. Naturally, the same is true for PNR and IS approaches. Furthermore, we show that the PUL parametrization and the Carroll compatible connection can be constructed by gauging the Carroll algebra, i.e., the equivalence of PUL and CAG. Similarly, we also prove that we can construct the PNR parametrization and the Galilei compatible connection from the Galilei algebra gauging procedure, i.e., the equivalence of PNR and GAG.
@@ II Review of non-Lorentzian limits
We begin with a brief review of the methods that are used in the available literature for construction of non-Lorentzian theories.
@@ II.1 Pre-ultralocal parametrization
The PUL parameterization is the parametrization that is suitable for taking the ultralocal (Carrollian) limit [64]. The metric parametrization is the given by
It proves useful to introduce a Carroll compatible connection by demanding that $\Pi^{\mu\nu}$ and $T_{\mu}$ are covariantly constant,
@@ II.2 Pre-nonrelativistic parametrization
A convenient parametrization of the metric that allows one to take the nonrelativistic (Galilean) limit is the PNR parametrization [68]. The metric parameterization is the same as (1) with the conditions (2), but a different connection that is compatible with Galilean symmetries:
@@
Now, assuming the analyticity of all the quantities, we expand $\Pi_{\mu\nu},\Pi^{\mu\nu},V^{\mu},T_{\mu}$ in powers of ${c^{-2}}$:
@@ II.3 Zero signature approach
The ZS approach is a method of taking the ultralocal limit of a Lorentzian theory. It consists of performing the ADM decomposition, and then setting the signature of the manifold to zero. This process results in a Carrollian manifold [79]. This is equivalent of taking the limit of the norm of the normal vector to zero. The first step is to decompose the metric with a Lorentzian and a Riemannian signatures. The Lorentzian metric is decomposed as
$g_{\mu\nu}=-n_{\mu}n_{\nu}+h_{\mu\nu},\hskip 10.0pt$ $\hskip 10.0ptg^{\mu\nu}=-n^{\mu}n^{\nu}+h^{\mu\nu},$ (7)
where $h_{\mu\nu}$ is the induced metric on the spatial surfaces and $n_{\mu}$ is the corresponding orthogonal 1-form with $n_{\mu}n^{\mu}=-1$. The Riemannian counterpart is
$g_{\mu\nu}=n_{\mu}n_{\nu}+h_{\mu\nu},\hskip 10.0pt$ $\hskip 10.0ptg^{\mu\nu}=n^{\mu}n^{\nu}+h^{\mu\nu}.$ (8)
Calculating the Hamiltonian of GR in both signatures yields
$H=G_{\mu\nu\sigma\rho}\pi^{\mu\nu}\pi^{\sigma\rho}-\sqrt{g}R=G_{\mu\nu\sigma\rho}\pi^{\mu\nu}\pi^{\sigma\rho}+\bar{R}-(K^{2}-K_{\mu\nu}K^{\mu\nu}),$ (9)
where $G_{\mu\nu\sigma\rho}=\tfrac{1}{2\sqrt{g}}(g_{\mu\sigma}g_{\nu\rho}+g_{\mu\rho}g_{\nu\sigma}-g_{\mu\nu}g_{\sigma\rho})$, $\bar{R}$ is the three dimensional Ricci scalar, $K_{\mu\nu}$ is the extrinsic curvature, and $K=h^{\mu\nu}K_{\mu\nu}$ in the Lorentzian case, and
$H=G_{\mu\nu\sigma\rho}\pi^{\mu\nu}\pi^{\sigma\rho}+\sqrt{g}R=G_{\mu\nu\sigma\rho}\pi^{\mu\nu}\pi^{\sigma\rho}-\bar{R}-(K^{2}-K_{\mu\nu}K^{\mu\nu}),$ (10)
in the Riemannian case. The first term vanishes by the Hamiltonian constraints and we end up with the same results as if we began with the Lagrangian.
It was shown in [79] that removing the terms that switches sign with signature results in a quantity that respects Carroll symmetries, i.e, $K^{2}-K_{\mu\nu}K^{\mu\nu}$. This is effectively the same as setting the signature of the manifold to zero. We will do the same procedure using the Lagrangian formalism in the following section in a different way, by considering the signature as a real parameter.
Contrary to the previously reviewed approaches, here we use the usual Levi-Civita connection. Expressing the Lagrangian in terms of the ADM variables, then setting the signature of the manifold to zero, we get a degenerate metric with the Carrollian theory (on a Carrollian manifold).
II.4 Infinite scaling approach
Starting from a Lorentzian theory and writing the metric in the same form as in the ZS approach but taking the limit of the norm of the normal vector to $\infty$- ∞ results in a Galilean theory. Alternatively, starting with the Riemannian theory and taking the limit to $\infty$ results in the same Galilean theory; the difference would be only in an overall sign in the Lagrangian.
II.5 Carroll algebra gauging
@@
The second equality follows from $\xi^{\mu}\omega_{\mu}^{a}=\xi^{\mu}\omega_{\mu\nu}^{ab}=0$. Let $\omega$ be a real parameter. We notice that the rescaling of $\tau_{\mu}$ and $\zeta$ by $\omega^{-1}$ and of $\omega_{\mu}{}^{a}$ and $\lambda^{a}$ by $\omega$ leaves the connection and the infinitesimal transformation parameter invariant. The field strength tensor (curvature) associated to $\mathcal{A}_{\mu}$ as
@@
II.6 Galilei algebra gauging
@@
where $G_{a}$ denote the Galilean boosts. The general connection and the infinitesimal transformation parameter are the same as in the Carrollian case. However, the rescaling of $\tau_{\mu}$ and $\zeta$ by $\omega$, and of $\omega_{\mu}{}^{a}$ and $\lambda^{a}$ by ${\omega}^{-1}$ leaves the connection and the infinitesimal transformation parameter invariant. The curvatures are given by
@@
where $\partial_{\mu}$ is the partial derivative with respect to a generic coordinate system. Taking the limit $\omega\to\infty$ gives Galilean theories.
@@ III.1 Equivalence of PUL and ZS approaches
Here, we rewrite the formulae used in the ADM decomposition with an extra real parameter $\epsilon$. This parameter should be understood as a modification to the normalization of the vector $n_{\mu}$ that is orthogonal to the spacelike slices, but not necessarily orthonormal, ${n_{\mu}n^{\mu}=\epsilon}$. The metric is Lorentzian for ${\epsilon<0}$, Riemannian for ${\epsilon>0}$, and degenerate for ${\epsilon=0}$. Following the usual ADM decomposition procedure, we write the metric and its inverse in an adapted coordinate system satisfying ${\boldsymbol{\partial}_{t}=N\boldsymbol{n}+\boldsymbol{N}}$,
@@
$g^{\mu\nu}=\begin{bmatrix}\tfrac{1}{\epsilon N^{2}}&\tfrac{N^{j}}{N^{2}}\\ \tfrac{N^{i}}{N^{2}}&h^{ij}+\epsilon\frac{N^{i}N^{j}}{N^{2}}\end{bmatrix},$ (20)
@@
One can show that $h_{\mu}^{\nu}h_{\nu}^{\rho}=h^{\rho}_{\mu}$, and $h^{\mu}_{\nu}=\delta_{\nu}^{\mu}-\tfrac{1}{\epsilon}n^{\mu}n_{\nu}$. The extrinsic curvature is then given by
$K_{\mu\nu}=h_{\mu}^{\alpha}h_{\nu}^{\beta}\nabla_{\alpha}n_{\beta}=-\nabla_{\nu}n_{\nu}+\tfrac{1}{\epsilon}n_{\mu}n^{\alpha}\nabla_{\alpha}n_{\nu}.$ (23)
It is easy to deduce that $\boldsymbol{K}=-\tfrac{1}{2}\pounds_{\boldsymbol{n}}\boldsymbol{h}$. We notice that although (21) was written in some adapted coordinates the fact that $n_{\mu}$ has an overall factor of $\epsilon$ while $n^{\mu}=O(1)$ as $\epsilon\to 0$ does not change. This is because the dependency on the signature does not depend on the choice of the coordinate system. Thus, we can rescale $n_{\mu}=-\epsilon\tilde{n}_{\mu}$ where $\tilde{n}_{\mu}=O(1)$ when $\epsilon\to 0$. Writing (22) in terms of the covector $\tilde{n}_{\mu}$ we get
@@
Having shown that, we write Gauss relation and its contractions in terms of the parameter $\epsilon$. Let us begin with the Gauss relation,
@@ III.2 Equivalence of PNR and IS approaches
Similar to the previous subsection, starting from (24) and taking the limit $\epsilon\to-\infty$, results in the metric in [68] in the PNR parameterization. Thus, all the curvature tensors in both approaches must be equivalent. All the formulae afterwards are valid for the Galilean limit provided (in terms of a different connection) we take the limit $\epsilon\to-\infty$. The limit $\epsilon\to\infty$ gives the same formulae except for an overall negative sign.
@@ III.3 Equivalence of PUL and CAG approaches
Let us now reconstruct the PUL parametrization from the CAG procedure. Using the procedure delineated in Sec. II.5, we define inverses to the gauge fields vielbein ($\tau_{\mu}$ and $e_{\mu}{}^{a}$) by
@@
Performing the redefinitions in [14], where $\tau_{\mu}$ , $\omega_{\mu}{}^{a}$, $\zeta$, and $\lambda^{a}$ are rescaled by $\omega^{-1}$ and taking the limit $\omega\to\infty$, leaving $\mathcal{A}_{\mu}$ and $\Pi$ invariant, we can write the metric and its inverse as
@@
This form is the same form in the PUL approach with ${\omega=c^{-2}}$. It follows from this fact is that all curvature tensors will be the same in the two approaches. Thus, any gravity theory will be expanded similarly, i.e, the two approaches are equivalent.
We now derive the Carroll compatible connection presented in [64] from the CAG approach. To do that, let us define a covariant derivative that are compatible with the veilbein [14] by
@@
(For more general connections see [62].) This procedure is the generalized version of the one in appendix B of [64]. It is easy to see from this point using (16) that the torsion and the curvature tensors match that of the PUL ones. For example, the Riemann tensor is given by
@@
which, upon calculating explicitly, gives the same form as (2.18) in [85].